OpenStudy (anonymous):

Which of the following is a solution of x2 + 4x = −8? x = −2 − 2i x = 2 + 2i x = −4 + 4i x = 4 − 4i

OpenStudy (anonymous):

helppp plzzz

OpenStudy (anonymous):

solving a quadratic equation is easy , you can do it by either splitting the middle term (in this case its not possible since the roots/ solution is not real) so you have to solve it by finding he discriminant write the equation in the form of \[ax ^{2}+bx+c=0\] where a is the coefficient of x^2, b is the coefficient of x and c is the constant term discriminant is given by : \[D=b ^{2}-4ac\] now to find the value of x we use the equation: \[x=(-b \pm \sqrt{D})/2a\]

OpenStudy (anonymous):

so its C

OpenStudy (anonymous):

no it would be A

OpenStudy (anonymous):

see D=(4)^2-4(1)(8) =16-32 =-16 so x=(-4-sqrt(D))/2(1) and (-4+sqrt(D))/2(1) =(-4-sqrt(-16))/2 and (-4+sqrt(-16))/2 =(-4-4i)/2 and (-4+4i)/2 so the two solutions are: -2-2i and -2+2i